NEW EXPERIMENTAL DATA OF THE WATER VAPOR ABOVE ∼ 25000 cm^{−1} BY CW-CRLAS

TITUS GHERMAN and \underline{PATRICK DUPRÉ}, Laboratoire de Spectrométrie Physique, Université Joseph Fourier, 140 Avenue de la Physique, BP 87, 38402 Saint Martin d’Hères, Cedex, France.

The \underline{C}avity \underline{R}ingDown \underline{L}aser \underline{A}bsorption \underline{S}pectroscopy (CRLAS or CRDS) technique is a unique tool for probing weakly absorbing molecular species. High resolution (Doppler-limited broadening) is achievable in the blue-UV energy region by intracavity frequency doubling a CW laser source (here a Ti:Sa laser) externally. Moreover, absolute intensity and energy calibrations are available. The H_{2}O molecule in vapor phase exhibits a weak absorption spectrum in the visible-UV energy range. Fourier transform spectra are available up to ∼ 25000 cm^{−1}; our experimental setup allows us to probe the UV energy range potentially up to ∼ 27000 cm^{−1} (i.e. the energy region of the polyad 8ν, 8ν + δ) with Doppler-limited linewidth and a sensitivity of ∼ 4 ppb/cm/\sqrt{Hz}. The modelling of high vibrational energy levels is now available as part of the ab initio investigation of floppy molecules. The new experimental data show poor agreement compared to the available ab initio predictions.